# BlockFusion: Expandable 3D Scene Generation using Latent Tri-plane Extrapolation

### v3

@@ Abstract.
To expand a scene during generation, one needs only to append empty blocks to overlap with the current scene and extrapolate existing latent tri-planes to populate new blocks. The extrapolation is done by conditioning the generation process with the feature samples from the overlapping tri-planes during the denoising iterations. Latent tri-plane extrapolation produces semantically and geometrically meaningful transitions that harmoniously blend with the existing scene. A 2D layout conditioning mechanism is used to control the placement and arrangement of scene elements. Experimental results indicate that BlockFusion is capable of generating diverse, geometrically consistent and unbounded large 3D scenes with unprecedented high-quality shapes in both indoor and outdoor scenarios. A demo video can be found at https://www.youtube.com/watch?v=PxIBtd6G0mA.
@@ 1. Introduction
Generating large amount of high-quality 3D contents is key for many practical applications, including video-games, film-making, augmented and virtual reality (AR/VR). The increasing demand for high-quality digital contents has made 3D generation a significant topic of research. Recently, in the 2D domain, denoise diffusion models (Sohl-Dickstein et al., 2015), have demonstrated remarkable results in image synthesis and beyond, leading to the development of production-ready 2D generation tools, such as Stable Diffusion (Rombach et al., 2022a), Midjourney, and Dall-E (Ramesh et al., 2021). The success in 2D domain has significantly sparked interest in the development of 3D generation tools. A multitude of researches on 3D generation have been published recently, most notable works include DreamFusion (Poole et al., 2022), Rodin (Wang et al., 2023b), Get3D (Gao et al., 2022), Zero123 (Liu et al., 2023d), SyncDreamer (Liu et al., 2023b), and LRM ((Hong et al., 2023)), etc.
@@
Instead of generating 3D through 2D image lifting, another research direction involves directly learning to produce 3D data, using supervision from either 3D shape ground truths or posed multi-view images. Notable methods include EG3D (Chan et al., 2022), Rodin (Wang et al., 2023b), and Get3D (Gao et al., 2022), etc. These approaches represent 3D data with a continuous hybrid neural field architecture, typically consisting of a tri-plane and an MLP decoder. The tri-plane, originally introduced in EG3D (Chan et al., 2022) and (Peng et al., 2020), is a tensor used to factorize the dense 3D volume grid. It is built on three axis-aligned 2D planes: the XY, YZ, and XZ planes. The MLP decoder converts the tri-plane feature into a continuous value representing the scene, which could be occupancy, signed distance field (SDF), radiance field (Mildenhall et al., 2021), etc. Tri-plane is significantly more compact and computationally efficient than a full 3D tensor and conducive to generative architectures developed for 2D image synthesis. This has been a key factor in making high-quality direct 3D data generation possible.
@@ 2.1. Diffusion models
The success of 2D generation tools based on diffusion models, notably Stable Diffusion(Rombach et al., 2022b), Midjourney, and Dall-E, has significantly sparked interest in the development of 3D generation tools. There are two main streams for this task: the methods that lift 2D (generated) images into 3D models, and the methods that directly run diffusion on 3D data. A thorough review of diffusion models for visual computing is available in (Po et al., 2023).
@@
3D diffusion models. Another line of research involves directly training diffusion models to generate 3D shapes. As the supervision comes directly from 3D shape ground truth or posed multi-view images, the generated results typically exhibit superior geometric quality compared to those from 2D diffusion-based methods. These methods can be categorized based on the type of 3D representations they employ, including: polygon meshes (Gao et al., 2022; Liu et al., 2023a), point clouds (Nichol et al., 2022; Zeng et al., 2022), explicit 3D grids holding occupancy or SDF values (Zheng et al., 2023; Liu et al., 2023c), or neural fields (Wang et al., 2023b; Shue et al., 2023; Chen et al., 2023a; Müller et al., 2023; Jun and Nichol, 2023; Xu et al., 2023b; Erkoç et al., 2023; Chou et al., 2023). The hybrid neural field, which incorporates a tri-plane followed by a neural decoder, has been widely adopted in 3D diffusion models due to its computational efficiency. Rodin (Wang et al., 2023b) first fits tri-plane NeRFs for a human upper body dataset, and then uses a two-stage coarse-to-fine diffusion model to generate the corresponding tri-planes. Similarly, NFD (Shue et al., 2023) trains a tri-plane diffusion model for 3D data parametrized via occupancy values. SSDNerf (Chen et al., 2023a) merge tri-plane fitting and generation into a single-stage pipeline. However, in practice, tri-plane diffusion is still challenging to train due to its high dimensionality and irregularity. Existing methods only demonstrated simple cases with small data varieties, i.e., Rodin for canonicalized human upper-body dataset (Wood et al., 2021), and NFD and SSDNeRF for single-category objects in ShapeNet (Chang et al., 2015). This paper follows this line of research but introduces a major change: we use an auto-encoder to compress the tri-plane into a highly compact latent tri-plane space for diffusion. We demonstrate that this approach significantly improves the stability, generalizability, and output quality of tri-plane diffusion.
@@
Generating 3D scenes presents a more substantial challenge than generating single objects. This is because scenes are geometrically more complex than individual objects, and they cannot be contained in a fixed spatial size. Object retrieval-based approaches assume there is a database of objects, and they arrange the retrieved objects to fill an empty scene, as seen in Diffuscene (Tang et al., 2023a) and Sceneformer (Wang et al., 2021), consequently, the synthesized scene can not contain novel elements that do not exist in the database. Text2Room (Höllein et al., 2023) is the first method that uses 2D diffusion model to build a 3D generation tool. It first generates color and depth frames using 2D diffusion models, and then shift camera positions to generate new frames, which are integrated into a global map. A similar approach for indoor scenarios can be found in SceneScape (Fridman et al., 2023). To allow precise control over the contents generated in a scene, ControlRoom3D(Schult et al., 2023) and CTRL-ROOM(Fang et al., 2023) develop panorama-based room generation models that take 3D room layouts as input conditions. CC3D (Bahmani et al., 2023) utilizes a 3D layout to guide the SDS (Poole et al., 2022) process. Citygen (Deng et al., 2023) represents city scenes using the height map proxy, leading to a 2.5D scene generation. Other approaches, such as (Tang et al., 2023b; Chen et al., 2023c; Wang et al., 2023a) focus on generating scenes with high-quality visual appearances. Given a room scene mesh, MVDiffusion (Tang et al., 2023b) generates coherent multiview perspective images, which can be lifted to the 3D as the UV texture of the mesh.
SceneDreamer (Chen et al., 2023c) use in-the-wild 2D image collections to learn a feed-forward generative model for unbounded 3D scenes. Semantic height maps are required while inference.
@@ 3.3. Compressing to latent tri-plane space
Although our raw tri-planes can reconstruct high-quality shapes, we found that generating such tri-planes is significantly difficult. Directly training diffusion models on such tri-planes leads to collapsed results, as shown in Fig. 4. We argue that there are mainly two reasons for this: 1) the raw tri-plane is highly redundant, and 2) the shape diversity in our scene block dataset is too large. Although previous works like Rodin (Wang et al., 2023b) and NFD (Shue et al., 2023) have proven the feasibility of diffusion on raw tri-planes, they only work on datasets with much smaller varieties, i.e., Rodin for canonicalized human upper bodies, and NFD for single-category objects from ShapeNet (Chang et al., 2015). When we attempted to retrain NFD on our scene blocks, it also failed to produce meaningful shapes, as shown in Fig. 4.
@@
where $\mathcal{L}_{rec}$ is a light $L_{1}$ norm applied between $x$ and its reconstruction $\mathcal{D}(\mathcal{E}(x))$. $\mathcal{L}_{KL}$ is a the Kullback-Leibler-term between $q_{\mathcal{E}}(z|x)=\mathcal{N}(z;\mathcal{E}_{u},\mathcal{E}_{\sigma^{2}})$ and a standard normal distribution $N(z;0,1)$ as in a standard vae (Kingma and Welling, 2013). To obtain high-fidelity shape reconstructions we only use a very small weight for $\mathcal{L}_{KL}$. $\mathcal{L}_{geo}$ is the geometry loss defined in Eqn. 2. It is assessed based on the same set of points as in Sec. 3. Since the purpose is to learn a latent tri-plane that can faithfully represent the shape, we rely on $L_{geo}$ as the dominate loss for training the auto-encoder.
@@ 3.4. Latent Triplane Diffusion
There are different ways to parameterize $\mu_{\Psi}(z_{t},t)$ in the prior. Instead of predicting the added noise as in the original DDPM, in this paper, we predict $z_{0}$ directly with a neural network $\Psi$. The prediction could be used in Eqn. 14 to produce $\mu_{\Psi}(z_{t},t)$. Specifically, with a uniformly sampled time step $t$ from $\{1,...,T\}$, we sample noise to obtain $z_{t}$ from input latent vector $z_{0}$. A time-conditioned denoising auto-encoder $\Psi$ learns to reconstruct $z_{0}$ from $z_{t}$. The objective of latent tri-plane diffusion reads
@@
where $\gamma(\cdot)$ is a positional encoding function and $||\cdot||_{2}$ is MSE loss. Since the forward process is fixed, $z_{t}$ can be efficiently obtained from $\mathcal{E}$ during training. During test time, we iteratively denoise $z_{T}$ until we obtain the final output $z^{\prime}$. $z^{\prime}$ can be decoded to the raw tri-plane $x^{\prime}$ with a single pass through $\mathcal{D}$. Finally, the pretrained MLP decodes $x^{\prime}$ to a dense SDF volume for marching cube-shape extraction.
@@
3D aware denoising U-Net. The neural backbone $\Psi(\cdot)$ of our model is realized as a time-conditional U-Net. The advantage of tri-planes is that we can treat them as 2D tensors and therefore apply efficient 2D convolutions. However, naively running convolution on tri-planes does not produce satisfactory results, as the 3D relationships among the plane features are ignored. $\Psi(\cdot)$ needs to incorporate operations that can account for the cross-plane feature relationships. To address this, Rodin (Wang et al., 2023b) introduces a 3D-aware convolution, which employs max-pooling and concatenation to associate features between planes based on their 3D correlations. However, the simple max-pooling max-pooling selects the largest one and discards the rest, inevitably causing information loss. In this work, we build $\Psi(\cdot)$ by leveraging the more powerful transformer to achieve cross-plane communication. The overall architecture of $\Psi(\cdot)$ is shown in Fig. 5. This architecture enables effective 3D-aware feature learning.
@@ 3.6. Surface refinement with non-rigid registration
Figure 9. Qualatitive room generation results. Text2Room (Höllein et al., 2023) generates distorted shapes and cannot accurately respond to the number of objects in the scene. For instance, when given the prompt ”one bed”, it generates multiple beds. In contrast, BlockFusion produces higher-quality shapes and correctly responds to numerical prompts.
@@ 4. Experimental Results
Datasets. We test our algorithm on three different types of scenes: room, city, and village. Room scene data is obtained from 3DFront (Fu et al., 2021a) and 3D-FUTURE (Fu et al., 2021b), which contains 18,968 indoor scenes with 34 classes of indoor objects. We obtain 57K random crops from 3DFront, with each block size set to $3.2^{3}$ cubic meters. We filtered out empty rooms and rooms with less than 5 objects and finally got 9123 rooms. For simplicity, we regroup the objects in 3DFUTURE based on their similarities into 9 classes: ”floor”, ”wall”, ”chair”, ”cabinet”, ”sofa”, ”table”, ”lighting”, ”bed”, and ”stool”. The city and village scenes are designed by artists, from each, we obtain 10K blocks. The block sizes are set to $12^{3}$ and $15^{3}$ cubic meters, respectively. The layout labels for the village are ”pine”, ”cypress tree”, ”ground”, and ”houses”, and for the city, they are ”road”, ”tree”, ”solar panels”, ”cars”, ”houses”. Note that all the blocks are cropped at random, and the testing blocks are never exposed to the model during training. During inference, the input semantic layout is created using an easy-to-use GUI, where user can place bounding boxes to indicate objects such as ”car” and ”tree”, or draw lines/contours to indicate continuous areas such as ”wall” and ”road”.
@@ 4.2. Evaluation Metrics
User study metric. We carried out a user study involving seven participants who were asked to rate the scene generation results based on Perceptual Quality (PQ) and Structure Completeness (SC) of the entire scene, using a scale from 1 to 5. This is done in two modes: textured mode (T-) and geometry-only mode (G-). In the Textured Mode, participants viewed a textured mesh, while in the Geometry-Only Mode, the texture was replaced with a monochrome material to emphasize the geometry. As a result, we derived four metrics from this study: T-PQ, T-SC, G-PQ and G-SC.
Table 1. Quantitative indoor scene generation results.
Textured Geometry-only
@@
Text2Room (Höllein et al., 2023) 2.14 2.29 1.00 1.28
Ours 4.14 4.14 3.71 3.86
@@
Single block generation. We regard NFD (Shue et al., 2023) as the baseline for single block generation. NFD is also based on tri-plane diffusion. However, they use occupancy values to represent shapes, whereas we employ SDF. We retrain NFD on our indoor scene blocks before evaluation. Tab. 2 and Fig. 4 show the results of unconditioned indoor block generation. Quantitatively, our method significantly outperforms NFD, with a $21.17\%$ and $23.67\%$ increase in coverage (Cov) scores under the CD and EMD metrics, respectively. Qualitatively, NFD is unable to generate meaningful shapes.
@@
Raw tri-plane Diff. (Ours) 0.0544 0.2744 23.99 27.50 89.91 88.50
Latent tri-plane Diff. (Ours) 0.0324 0.1884 51.83 53.33 70.66 60.08
Indoor scene generation. We consider Text2Room (Höllein et al., 2023) as the baseline for indoor scene generation. Text2Room takes text prompt as input whereas ours is based on 2D layout map. For a fair comparison, we describe our input room layout using natural language and then concatenate it as part of the text prompt for Text2Room. While our method does not directly generate textured mesh, we leverage an off-the-shelf text-to-texture generation tool, Meshy  [^1] https://www.meshy.ai/, to produce textures for our mesh. Meshy utilizes the same text prompt as Text2Room. To enhance the texture generation results of Meshy, we combine all blocks into a single entity using Blender’s voxel remeshing tool. Tab. 1 and Fig. 9 show the results of room generation. Qualitatively, due to the use of monocular depth estimation, the shape of Text2Room appears distorted, while BlockFusion produces significantly better room shapes. In addition, Text2Room cannot precisely react to the text prompt; it generates duplicate beds, while the prompt is ”one bed”. By leveraging layout control, our method can precisely determine the number of beds in the room. By leveraging Meshy, our approach also produces textures comparable to Text2Room. Quantitatively, questionnaires vote higher for BlockFusion.

### v4

@@ Abstract.
To expand a scene during generation, one needs only to append empty blocks to overlap with the current scene and extrapolate existing latent tri-planes to populate new blocks. The extrapolation is done by conditioning the generation process with the feature samples from the overlapping tri-planes during the denoising iterations. Latent tri-plane extrapolation produces semantically and geometrically meaningful transitions that harmoniously blend with the existing scene. A 2D layout conditioning mechanism is used to control the placement and arrangement of scene elements. Experimental results indicate that BlockFusion is capable of generating diverse, geometrically consistent and unbounded large 3D scenes with unprecedented high-quality shapes in both indoor and outdoor scenarios.
@@ 1. Introduction
Generating large amount of high-quality 3D contents is key for many practical applications, including video-games, film-making, augmented and virtual reality (AR/VR). The increasing demand for high-quality digital contents has made 3D generation a significant topic of research. Recently, in the 2D domain, denoise diffusion models (Sohl-Dickstein et al., 2015), have demonstrated remarkable results in image synthesis and beyond, leading to the development of production-ready 2D generation tools, such as Stable Diffusion (Rombach et al., 2022a), Midjourney, and Dall-E (Ramesh et al., 2021). The success in 2D domain has significantly sparked interest in the development of 3D generation tools. A multitude of researches on 3D generation have been published recently, most notable works include DreamFusion (Poole et al., 2022), Rodin (Wang et al., 2023), Get3D (Gao et al., 2022), Zero123 (Liu et al., 2023d), SyncDreamer (Liu et al., 2023b), and LRM ((Hong et al., 2023)), etc.
@@
Instead of generating 3D through 2D image lifting, another research direction involves directly learning to produce 3D data, using supervision from either 3D shape ground truths or posed multi-view images. Notable methods include EG3D (Chan et al., 2022), Rodin (Wang et al., 2023), and Get3D (Gao et al., 2022), etc. These approaches represent 3D data with a continuous hybrid neural field architecture, typically consisting of a tri-plane and an MLP decoder. The tri-plane, originally introduced in EG3D (Chan et al., 2022) and (Peng et al., 2020), is a tensor used to factorize the dense 3D volume grid. It is built on three axis-aligned 2D planes: the XY, YZ, and XZ planes. The MLP decoder converts the tri-plane feature into a continuous value representing the scene, which could be occupancy, signed distance field (SDF), radiance field (Mildenhall et al., 2021), etc. Tri-plane is significantly more compact and computationally efficient than a full 3D tensor and conducive to generative architectures developed for 2D image synthesis. This has been a key factor in making high-quality direct 3D data generation possible.
@@ 2.1. Diffusion models
The success of 2D generation tools based on diffusion models, notably Stable Diffusion(Rombach et al., 2022b), Midjourney, and Dall-E(Ramesh et al., 2021), has significantly sparked interest in the development of 3D generation tools. There are two main streams for this task: the methods that lift 2D (generated) images into 3D models, and the methods that directly run diffusion on 3D data. A thorough review of diffusion models for visual computing is available in (Po et al., 2023).
@@
3D diffusion models. Another line of research involves directly training diffusion models to generate 3D shapes. As the supervision comes directly from 3D shape ground truth or posed multi-view images, the generated results typically exhibit superior geometric quality compared to those from 2D diffusion-based methods. These methods can be categorized based on the type of 3D representations they employ, including: polygon meshes (Gao et al., 2022; Liu et al., 2023a), point clouds (Nichol et al., 2022; Zeng et al., 2022), explicit 3D grids holding occupancy or SDF values (Zheng et al., 2023; Liu et al., 2023c), or neural fields (Wang et al., 2023; Shue et al., 2023; Chen et al., 2023a; Müller et al., 2023; Jun and Nichol, 2023; Xu et al., 2023b; Erkoç et al., 2023; Chou et al., 2023). The hybrid neural field, which incorporates a tri-plane followed by a neural decoder, has been widely adopted in 3D diffusion models due to its computational efficiency. Rodin (Wang et al., 2023) first fits tri-plane NeRFs for a human upper body dataset, and then uses a two-stage coarse-to-fine diffusion model to generate the corresponding tri-planes. Similarly, NFD (Shue et al., 2023) trains a tri-plane diffusion model for 3D data parametrized via occupancy values. SSDNerf (Chen et al., 2023a) merge tri-plane fitting and generation into a single-stage pipeline. However, in practice, tri-plane diffusion is still challenging to train due to its high dimensionality and irregularity. Existing methods only demonstrated simple cases with small data varieties, i.e., Rodin for canonicalized human upper-body dataset (Wood et al., 2021), and NFD and SSDNeRF for single-category objects in ShapeNet (Chang et al., 2015). This paper follows this line of research but introduces a major change: we use an auto-encoder to compress the tri-plane into a highly compact latent tri-plane space for diffusion. We demonstrate that this approach significantly improves the stability, generalizability, and output quality of tri-plane diffusion.
@@
Generating 3D scenes presents a more substantial challenge than generating single objects. This is because scenes are geometrically more complex than individual objects, and they cannot be contained in a fixed spatial size. Object retrieval-based approaches assume there is a database of objects, and they arrange the retrieved objects to fill an empty scene, as seen in Diffuscene (Tang et al., 2023a) and Sceneformer (Wang et al., 2021), consequently, the synthesized scene can not contain novel elements that do not exist in the database. Text2Room (Höllein et al., 2023) is the first method that uses 2D diffusion model to build a 3D generation tool. It first generates color and depth frames using 2D diffusion models, and then shift camera positions to generate new frames, which are integrated into a global map. A similar approach for indoor scenarios can be found in SceneScape (Fridman et al., 2023). To allow precise control over the contents generated in a scene, ControlRoom3D(Schult et al., 2023) and CTRL-ROOM(Fang et al., 2023) develop panorama-based room generation models that take 3D room layouts as input conditions. CC3D (Bahmani et al., 2023) utilizes a 3D layout to guide the SDS (Poole et al., 2022) process. Citygen (Deng et al., 2023) represents city scenes using the height map proxy, leading to a 2.5D scene generation. PERF(Wang et al., 2024) proposes a novel panoramic view synthesis framework, thereby lifting 2D panorama up to 3D scene. Other approaches, such as (Tang et al., 2023b; Chen et al., 2023c) focus on generating scenes with high-quality visual appearances. Given a room scene mesh, MVDiffusion (Tang et al., 2023b) generates coherent multiview perspective images, which can be lifted to the 3D as the UV texture of the mesh. SceneDreamer (Chen et al., 2023c) leverages in-the-wild 2D images to construct large scenes with photo-realistic volume rendering effects. However, it still depends on 3D shapes represented by semantic height maps as input, consequently, the dimensions of the generated scenes are bounded by the dimensions of the input shapes.
@@ 3.3. Compressing to latent tri-plane space
Although our raw tri-planes can reconstruct high-quality shapes, we found that generating such tri-planes is significantly difficult. Directly training diffusion models on such tri-planes leads to collapsed results, as shown in Fig. 4. We argue that there are mainly two reasons for this: 1) the raw tri-plane is highly redundant, and 2) the shape diversity in our scene block dataset is too large. Although previous works like Rodin (Wang et al., 2023) and NFD (Shue et al., 2023) have proven the feasibility of diffusion on raw tri-planes, they only work on datasets with much smaller varieties, i.e., Rodin for canonicalized human upper bodies, and NFD for single-category objects from ShapeNet (Chang et al., 2015). When we attempted to retrain NFD on our scene blocks, it also failed to produce meaningful shapes, as shown in Fig. 4.
@@
where $\mathcal{L}_{rec}$ is a light $L_{1}$ norm applied between $x$ and its reconstruction $\mathcal{D}(\mathcal{E}(x))$. $\mathcal{L}_{KL}$ is a the Kullback-Leibler-term between $q_{\mathcal{E}}(z|x)=\mathcal{N}(z;\mathcal{E}_{u},\mathcal{E}_{\sigma^{2}})$ and a standard normal distribution $\mathcal{N}(\textbf{0},\textbf{I})$ as in a standard VAE (Kingma and Welling, 2013). To obtain high-fidelity shape reconstructions we only use a very small weight for $\mathcal{L}_{KL}$. $\mathcal{L}_{geo}$ is the geometry loss defined in Eqn. 2. It is assessed based on the same set of points as in Sec. 3. Since the purpose is to learn a latent tri-plane that can faithfully represent the shape, we rely on $L_{geo}$ as the dominate loss for training the auto-encoder. Detailed structures of VAE are presented in appendix.
@@ 3.4. Latent Triplane Diffusion
There are different ways to parameterize $\mu_{\Psi}(z_{t},t)$ in the prior. In this paper, we predict $z_{0}$ directly with a neural network $\Psi$. The prediction could be used in Eqn. 14 to produce $\mu_{\Psi}(z_{t},t)$. Specifically, with a uniformly sampled time step $t$ from $\{1,...,T\}$, we sample noise to obtain $z_{t}$ from input latent vector $z_{0}$. A time-conditioned denoising auto-encoder $\Psi$ learns to reconstruct $z_{0}$ from $z_{t}$. The objective of latent tri-plane diffusion reads
@@
where $\gamma(\cdot)$ is a positional encoding function and $||\cdot||_{2}$ is MSE loss. Since the forward process is fixed, $z_{t}$ can be efficiently obtained from $\mathcal{E}$ during training. During test time, we iteratively denoise $z_{T}$ until we obtain the final output $z^{\prime}$. $z^{\prime}$ can be decoded to the raw tri-plane $x^{\prime}$ with a single pass through $\mathcal{D}$. Finally, the pretrained MLP decodes $x^{\prime}$ to a dense SDF volume for shape extraction through marching cube.
@@
3D aware denoising U-Net. The neural backbone $\Psi(\cdot)$ of our model is realized as a time-conditional U-Net. The advantage of tri-planes is that we can treat them as 2D tensors and therefore apply efficient 2D convolutions. However, naively running convolution on tri-planes does not produce satisfactory results, as the 3D relationships among the plane features are ignored. $\Psi(\cdot)$ needs to incorporate operations that can account for the cross-plane feature relationships. To address this, Rodin (Wang et al., 2023) introduces a 3D-aware convolution, which employs max-pooling and concatenation to associate features between planes based on their 3D correlations. However, the simple max-pooling max-pooling selects the largest one and discards the rest, inevitably causing information loss. In this work, we build $\Psi(\cdot)$ by leveraging the more powerful transformer to achieve cross-plane communication. The overall architecture of $\Psi(\cdot)$ is shown in Fig. 5. This architecture enables effective 3D-aware feature learning.
@@ 3.6. Surface refinement with non-rigid registration
Figure 9. Qualitative room generation results. Text2Room (Höllein et al., 2023) generates distorted shapes and cannot accurately respond to the number of objects in the scene. For instance, when given the prompt ”one bed”, it generates multiple beds. In contrast, BlockFusion produces higher-quality shapes and correctly responds to numerical prompts.
@@ 4. Experimental Results
Datasets. We test our algorithm on three different types of scenes: room, city, and village. Room scene data is obtained from 3DFront (Fu et al., 2021a) and 3D-FUTURE (Fu et al., 2021b), which contains 18,968 indoor scenes with 34 classes of indoor objects. We obtain 57K random crops from 3DFront, with each block size set to $3.2^{3}$ cubic meters. We filtered out empty rooms and rooms with less than 5 objects and finally got 9123 rooms. For simplicity, we regroup the objects in 3D-FUTURE based on their similarities into 9 classes: ”floor”, ”wall”, ”chair”, ”cabinet”, ”sofa”, ”table”, ”lighting”, ”bed”, and ”stool”. The city and village scenes are designed by artists, from each, we obtain 10K blocks. The block sizes are set to $12^{3}$ and $15^{3}$ cubic meters, respectively. The layout labels for the village are ”pine”, ”cypress tree”, ”ground”, and ”houses”, and for the city, they are ”road”, ”tree”, ”solar panels”, ”cars”, ”houses”. Note that all the blocks are cropped at random, and the testing blocks are never exposed to the model during training. During inference, the input semantic layout is created using an easy-to-use GUI, where user can place bounding boxes to indicate objects such as ”car” and ”tree”, or draw lines/contours to indicate continuous areas such as ”wall” and ”road”.
@@ 4.2. Evaluation Metrics
User study metric. We carried out a user study involving 48 participants who were asked to rate the scene generation results based on Perceptual Quality (PQ) and Structure Completeness (SC) of the entire scene, using a scale from 1 to 5. This is done in two modes: textured mode (T-) and geometry-only mode (G-). In the Textured Mode, participants viewed a textured mesh, while in the Geometry-Only Mode, the texture was replaced with a monochrome material to emphasize the geometry. As a result, we derived four metrics from this study: T-PQ, T-SC, G-PQ and G-SC. The results are presented in Table 1 and 2.
Table 1. Quantitative indoor scene generation results of pure shape.
@@
Text2Room (Höllein et al., 2023) 1.92 1.92
Ours 4.44 4.58
Table 2. Quantitative indoor scene generation results of textured shape.
[…]
Text2Room (Höllein et al., 2023) 2.27 2.29
Ours + Meshy 4.08 4.17
@@
Single block generation. We regard NFD (Shue et al., 2023) as the baseline for single block generation. NFD is also based on tri-plane diffusion. However, they use occupancy values to represent shapes, whereas we employ SDF. We retrain NFD on our indoor scene blocks before evaluation. Tab. 3 and Fig. 4 show the results of unconditioned indoor block generation. Quantitatively, our method ($z_{0}$ target) significantly outperforms NFD, with a $29.17\%$ and $23.67\%$ increase in coverage (Cov) scores under the CD and EMD metrics, respectively. Qualitatively, NFD is unable to generate meaningful shapes.
@@
Raw tri-plane Diff. w/ $z_{0}$ (Ours) 0.0544 0.2744 23.99 27.50 89.91 88.50
Latent tri-plane Diff. w/ $z_{0}$ (Ours) 0.0324 0.1884 51.83 53.33 70.66 60.08
Latent tri-plane Diff. w/ $noise$ (Ours) 0.0326 0.1865 50.49 55.66 69.74 56.66
Indoor scene generation. We consider Text2Room (Höllein et al., 2023) as the baseline for indoor scene generation. Text2Room takes text prompt as input whereas ours is based on 2D layout map. For a fair comparison, we describe our input room layout using natural language and then concatenate it as part of the text prompt for Text2Room. Since our method does not directly generate textured meshes, we leverage an off-the-shelf text-to-texture generation tool, Meshy  [^1] https://www.meshy.ai/, to produce textures for our mesh. Meshy utilizes the same text prompt as Text2Room. To enhance the texture generation results of Meshy, we combine all blocks into a single entity using Blender’s voxel remeshing tool. Tab. 1, Tab. 2 and Fig. 9 show the results of room generation. Qualitatively, due to the use of monocular depth estimation, the shape of Text2Room appears distorted, while BlockFusion produces significantly better room shapes. In addition, Text2Room cannot precisely react to the text prompt; it generates duplicate beds, while the prompt is ”one bed”. By leveraging layout control, our method can precisely determine the number of beds in the room. By leveraging Meshy, our approach also produces textures comparable to Text2Room. Quantitatively, under a five-point system, Blockfusion are leading by $2.52$ points in geometric perceptual quality and $2.66$ points in geometric structure completeness respectively. In the case of textured generation, Blockfusion are leading by $1.81$ points in textured perceptual quality and $1.88$ points in textured structure completeness respectively.